OpenStudy (anonymous):

15. x+y+z-6=0 2x-3y+4z-3=0 4x-8y+4z-12=0 17.w+x+y+z=0 2w+x-y-z=1 -w-x+y+z=o 2x+y=o

OpenStudy (amistre64):

1 1 1 = 6 2 -3 4 = 3 4 -8 4 = 12 1 1 1 = 6 (-2) -2 -2 -2 =-12 2 -3 4 = 3 2 -3 4 = 3 ------------- 0 -5 2 = -9 .......................................................... 1 1 1 = 6 (*-4) -4 -4 -4 = -24 4 -8 4 = 12 4 -8 4 = 12 -------------- 0 -12 0 = -12 (/-12) 0 1 0 = 1 ......................................................... 0 -5 2 = -9 0 -5 2 = -9 0 1 0 = 1 (5) 0 5 0 = 5 -------------- 0 0 2 = -4 (/2) 0 0 1 = -2

OpenStudy (amistre64):

1 1 1 = 6 0 1 0 = 1 0 0 1 = -2 1 1 -2 = 6 1 0 0 = 6-1+2 x y z =# --------- 1 0 0 = 7 0 1 0 = 1 0 0 1 = -2

OpenStudy (amistre64):

1 1 1 1 = 0 2 1 -1 -1 = 1 -1 -1 1 1 = 0 2 1 0 0 = 0 1 1 1 1 = 0 -1 -1 1 1 = 0 (add em up, they are good to go) --------------- 0 0 2 2 = 0 .............................................................. 1 1 1 1 = 0 (*-2) -2 -2 -2 -2 = 0 2 1 -1 -1 = 1 2 1 -1 -1 = 1 --------------- 0 -1 0 -3 = 1 .................................................................. 1 1 1 1 = 0 (*-2) -2 -2 -2 -2 = 0 2 1 0 0 = 0 2 1 0 0 = 0 --------------- 0 -1 -2 -2 = 0 ............................................................... 1 1 1 1 = 0 0 -1 0 -3 = 1 (*-1) 0 1 0 3 = -1 0 -1 0 -3 = 1 0 -1 -2 -2 = 0 ............................................................... 0 1 0 3 = -1 0 -1 0 -3 = 1 -------------- 0 0 0 0 = 0 this is a sign for infinity solutions ........................................... 0 1 0 3 = -1 0 -1 -2 -2 = 0 ----------------- 0 0 -2 1 = -1

OpenStudy (amistre64):

1 1 1 1 = 0 0 1 0 3 = -1 0 0 -2 1 = -1 (/-2) 0 0 1 -1/2 = 1/2 0 0 0 0 = 0 ................................................ 1 1 1 1 = 0 0 1 0 3 = -1 0 0 1 -1/2 = 1/2 0 0 0 0 = 0

OpenStudy (amistre64):

w+x+y+z=0 2w+x-y-z=1 -w-x+y+z=0 y = -2x ------------------- w-x+z=0 w = x-z 2(x-z)+3x-z=1 -- 2x-2z+3x-z = 1 -(x-z)-3x+z=0 -- -x +z -3x+z = 0 ..................................... 5x -3z = 1 -4x +2z = 0 -- z = 2x 5x -2x = 1 3x = 1 x = 1/3 z = 2(1/3) = 2/3 w = x-z = 1/3 - 2/3 = -1/3 y = -2x = -2(1/3) = -2/3

OpenStudy (amistre64):

i see the typo in my matrix that goofed me up lol.... -2 -2 -2 -2 = 0 2 1 -1 -1 = 1 --------------- 0 -1 -3 -2 = 1 ^^ i made that a '0' last time :)

OpenStudy (anonymous):

was about to say the book says answer is 1, -1, 2, -2

OpenStudy (anonymous):

shld i erase up to the mistake?

OpenStudy (amistre64):

you could erase it up to that point :)

OpenStudy (anonymous):

got it to 1 0 0 0 1 0 1 00 -1 0 0 1 1 0 0 -1 -2 -2 0

OpenStudy (anonymous):

YES ALL THE STUFF U TAUGHT ME MAKES SENSE HAHAHA YES 1 SEC WANNA SEE IF I CAN GET TO 1 0 0 0 0 1 0 0 0 0 1 0 0 0 0 1 0

OpenStudy (anonymous):

sry caps

OpenStudy (amistre64):

lol....little excited eh :)

OpenStudy (anonymous):

well i was then i forgot wtf went thru my brain can get to 1 0 0 0 1 0 1 0 0 -1 0 0 1 1 0 0 -1 -2 -2 0

OpenStudy (anonymous):

what shld i do now?

OpenStudy (amistre64):

0 0 1 1 = 0 0 -1 -2 -2 = 0 switch these 2 rows 1 0 0 0 = 1 0 1 0 0 = -1 0 -1 -2 -2 = 0 (*1) 0 0 1 1 = 0

OpenStudy (amistre64):

or subtract it from the R2

OpenStudy (anonymous):

oh didnt know i could swap rows

OpenStudy (amistre64):

0 1 0 0 =-1 0-1-2-2= 0 ------------- 0 0 2 2 =-1 (/2) 0 0 1 1 = -1/2 .....maybe?

OpenStudy (amistre64):

matrix can be cumbersome.....

OpenStudy (anonymous):

nah cause the answers are 1, -1 2, -2

OpenStudy (anonymous):

and thats and understatement

OpenStudy (amistre64):

-2 -2 -2 -2 = 0 2 1 -1 -1 = 1 --------------- 0 -1 -3 -3 = 1 (*1) 0 1 3 3 = 1 ..............................

OpenStudy (anonymous):

wait how does that work cause if you times whole line by -1 to make them positives it makes the 1 -1 right?

OpenStudy (amistre64):

1 1 1 1 = 0 0 1 3 3 = 1 0 1 3 3 = 1 0 -1 -2 -2 = 0 0 -1 -2 -2 = 0 0 0 2 2 = 0 -------------- 0 0 -1 -1 =1 (-1) 0 0 1 1 = -1 .......................................................... 0 1 3 3 = 1 0 0 2 2 = 0 ------------- 0 1 5 5 = 1

OpenStudy (anonymous):

wait im confused if ur adding 0 1 3 3 1 to 0 -1 -2 -2 0 wouldnt it be 0 0 1 1 1? then when times by -1 0 0 -1 -1 -1?

OpenStudy (amistre64):

1 1 1 1 = 0 0 1 3 3 = 1 0 0 1 1 = -1 ...cant seem to wrap this up lol

OpenStudy (amistre64):

im gonna write this on paper; its a nightmare trying to keep things sorted in this reply box lol

OpenStudy (anonymous):

OMFG WE FAILED SO HARD AHHAHAHHAHA

OpenStudy (anonymous):

4rth row is not 2 1 0 0 y is third in all the others so wouldnt it be 2 0 1 0

OpenStudy (anonymous):

sry 0 2 1 0

OpenStudy (amistre64):

lol...your right

OpenStudy (anonymous):

always goes w x y z we have x and y

OpenStudy (anonymous):

haha if only i could fit all of this onto failblog or unfreindable

OpenStudy (amistre64):

1 1 1 1 =0 2 1 -1-1 =1 -1 -1 1 1 =0 0 2 1 0 =0 lol yeah :)

OpenStudy (anonymous):

dear god getting a drink and cranking this out hey btw meant to ask you earleir are you a teacher or what do you do?

OpenStudy (amistre64):

im an unemployed carpenter going back to college :)

OpenStudy (anonymous):

well congratz on going back my brother is going back soon too and my other brother is just about to finish at the community but im going into the navy already signed why i have to get this stuff done

OpenStudy (amistre64):

1 1 11=0 -1-111=0 ---------- 0 0 2 2 = 0 -2-2-2-2=0 2 1 -1-1=1 ------------ 0-1-3-3=1 -2-2-2-2=0 0 2 1 0 =0 ------------ -2 0-1-2=0

OpenStudy (amistre64):

1 1 1 1 = 0 0 1 3 3 =-1 0 0 1 1 =0 -2 0-1-2=0

OpenStudy (anonymous):

R1 times 2 add to R4 and swap R4 and R3?

OpenStudy (amistre64):

pawn to queen3 lol

OpenStudy (anonymous):

rofl nvm

OpenStudy (amistre64):

2 2 2 2=0 -20-1-2=0 ----------- 0 2 1 0 = 0 ....cant hurt lol

OpenStudy (anonymous):

oh times R3 by -3 add to R2?

OpenStudy (anonymous):

hahah like an undobutton

OpenStudy (amistre64):

0 1 3 3 =-1 (*-2) 0 2 1 0 = 0 0-2-6-6 =2 0 2 1 0 =0 ----------- 0 0 -5 -6 = 2

OpenStudy (amistre64):

0 0 1 1 =0 (*5) 0 0 -5 -6 = 2 00 5 5=0 00 -5-6 =2 ----------- 0 0 0 -1 = 2 0001=-2

OpenStudy (amistre64):

1 1 1 1 = 0 0 1 3 3 =-1 0 0 1 1 = 0 1 -2 = 0; 1 0=2 0 0 0 1 =-2

OpenStudy (amistre64):

1 1 1 1 = 0 0 1 3 3 =-1 <-- 0 1 6 -6 = -1 ; 0100=-1 0 0 1 0=2 0 0 0 1 =-2

OpenStudy (amistre64):

1 1 1 1 = 0 <- 1 -1 2-2=0; 1000=1 2-2; 1000=1 0 1 0 0 =-1 0 0 1 0 = 2 0 0 0 1 =-2

OpenStudy (amistre64):

1 0 0 0 = 1 0 1 0 0 =-1 0 0 1 0 = 2 0 0 0 1 =-2

OpenStudy (amistre64):

did you get all that lol

OpenStudy (anonymous):

yes thx man for everything is there a way to add you so i cld maybe get future help if need be and ur ok with it?

OpenStudy (amistre64):

add me? not that i know of; just come by and if im here say hi :)

OpenStudy (amistre64):

do i deserve a medal :) lol

OpenStudy (anonymous):

i gave you one long time ago

OpenStudy (anonymous):

oh sweet can give you another

OpenStudy (amistre64):

true ;)

OpenStudy (amistre64):

yay lol..... thanx :)

OpenStudy (anonymous):

nah man thank you for everything i beat my head agianst this and i just couldnt make it come together in a way that made sense you showed me how to do it even tho my PAID teachers cldnt i honestly cld have failed if not for this then there wld be no navy nuclear school or graduation

OpenStudy (amistre64):

good luck :)